OpenStudy (anonymous):

solve the equation in the real numbers system: x^4-8x^3+16x^2+8x-17=0

myininaya (myininaya):

x=1 looks like it might work by inspection 1^4-8*1^3+16*1^2+8(1)-17 =1-8+16+8-17=1+16-17=0 so use syntehic division to find another factor besides (x-1)

myininaya (myininaya):

1 | 1 -8 16 8 -17 | 1 -7 9 17 -------------------------- 1 -7 9 17 | 0 x^3-7x^2+9x+17=0

OpenStudy (anonymous):

if it had another rational root it would have to be plus or minus 17. (Rational Root theorem)

myininaya (myininaya):

-1| 1 -7 9 17 | -1 8 -17 ----------------- 1 -8 17| 0

OpenStudy (anonymous):

or -1 as well, i always forget that >.<

myininaya (myininaya):

x=-1 works also

myininaya (myininaya):

so we have x^2-8x+17=0 we can solve this by using quadratic formula :) and then we are done

OpenStudy (anonymous):

1 | 1 -8 16 8 -17 | 1 -7 9 17 -------------------------- 1 -7 9 17 | 0 x^3-7x^2+9x+17=0 i dont really understand about this:(.

OpenStudy (anonymous):

how do u got that???

myininaya (myininaya):

synetic division?

myininaya (myininaya):

synthetic*

myininaya (myininaya):

if you want you can also do long division

OpenStudy (anonymous):

i dont know with the long division.. ok i'm understanding with the synthetic division now:)